# A Trio Neural Model for Dynamic Entity Relatedness Ranking

### v4

@@ 1 Introduction
  • We present the first study of dynamic entity relatedness ranking using collective attention.
  • We introduce an attention-based convolutional neural networks (CNN) to capture the temporal signals of an entity.
  • We propose a joint framework to incorporate multiple views of the entities, both from content provider and from user’s perspectives, for entity relatedness ranking.
@@ 3.2 Problem Definition
In our setting, it is not required to have a pre-defined, static function quantifying the semantic relatedness between two entities. Instead, it can capture a family of functions $F$ where the prior distribution relies on time parameter. We formalize the concepts below.
Dynamic Entity Relatedness between two entities $e_{s},e_{t}$, where $e_{s}$ is the source entity and $e_{t}$ is the target entity, in a given time $t$, is a function (denoted by $\mathsf{f_{t}}(e_{s},e_{t})$) with the following properties.
  • asymmetric: $f_{t}(e_{i},e_{j})\neq f_{t}(e_{j},e_{i})$
  • non-negativity: $f(e_{i},e_{j})\geq 0$
  • indiscernibility of identicals: $e_{i}=e_{j}\rightarrow f(e_{i},e_{j})=1$
Dynamic Entity Relatedness Ranking. Given a source entity $e_{s}$ and time point $t$, rank the candidate entities $e_{t}$’s by their semantic relatedness.
@@ 4.2 Neural Ranking Model Overview
The entity relatedness ranking can be handled by a point-wise ranking model that learns to predict relatedness score directly. However, as the navigational frequency distribution is often skewed at top, supervisions guided by long-tail navigations would be prone to errors. Hence instead of learning explicitly a calibrated scoring function, we opt for a pair-wise ranking approach. When applying to ranking top-$k$ entities, this approach has the advantage of correctly predicting partial orders of different relatedness functions $f_{t}$ at any time points regardless of their non-transitivity Cheng et al. (2012).
This work builds upon the idea of interaction-based deep neural models, i.e. learning soft semantic matches from the source-target entity pairs. Note that, we do not aim for a Siamese architecture Chopra et al. (2005) (i.e., in representation-based models), where the weight parameters are shared across networks. The reason is that, the conventional kind of network produces a symmetric relation, violating the asymmetric property of the relatedness function $f_{t}$ (section 3.2). Concretely, each deep network $\psi$ consists of an input layer $z_{0}$, $n-1$ hidden layers and an output layer $z_{n}$. Each hidden layer $z_{i}$ is a fully-connected network that computes the transformation: $z_{i}=\sigma(\mathbf{w}_{i}\cdot z_{i-1}+b_{i})$, where $\mathbf{w}_{i}$ and $b_{i}$ are the weight matrix and bias at hidden layer i, $\sigma$ is a non-linear function such as the rectified linear unit(ReLU). The final score under the trio setup is summed from multiple networks.
$\phi(<e_{s},e_{(+)},e_{(-)}>)=\phi_{time}+\phi_{graph}+\phi_{content}$ (1)
In the next section we describe the input representations $z_{0}$ for each network.
@@ 5.1 Content-based representation learning
To learn the entity representation from its content, we rely on entity textual document (word-based) as well as its link profile (entity-based) (section 3.1). Since the vocabulary size of entities and words is often very large, conventional one-hot vector representation becomes expensive. Hence, we adopt the word hashing technique from Huang et al. (2013), that breaks a term into character trigraphs and thus can dramatically reduce the size of the vector dimensionality. We then rely on embeddings to learn the distributed representations and build up the soft semantic interactions via input concatenation. Let $\mathsf{E}:\mathcal{V}\rightarrow\mathbb{R}^{m}$ be the embedding function, $\mathcal{V}$ is the vocabulary and $m$ is the embedding size. $\mathsf{w}:\mathcal{V}\rightarrow\mathbb{R}$, is the weighting function that learns the global term importance and a weighted element-wise sum of word embedding vectors -compositionality function $\oplus$, the word-based representation for entity e is hence $\oplus_{i=1}^{|e_{w}|}(\mathsf{E}(w_{i}),\mathsf{w}(w_{i}))$. For entity-based representation, we break down the surface form of a linked entity into bag-of-words and apply analogously. The concatenation of the two representations for the tuple $<e_{s},e_{(+)},e_{(-)}>$ is then input to the deep feed-forward network.
@@ 5.2 Graph-based representation
To obtain the graph embedding for each entity, we adopt the idea of DeepWalk Perozzi et al. (2014), which learns the embedding by predicting the vertex sequence generated by random walk. Concretely, given an entity $e$, we learn to predict the sequence of entity references $\mathbb{S}_{e}$ – which can be considered as the graph-wise context in the Skip-gram model. We then adopt the matching histogram mapping in Guo et al. (2016) for the soft interaction of the ranking model. Specifically, denote the bag of entities representation of $e_{s}$ as $\mathbb{C}_{e_{s}}$, and that of $e_{t}$ as $\mathbb{C}_{e_{t}}$; we discretize the soft matching (calculated by cosine similarity of the embedding vectors) of each entity pair in $(\mathbb{C}_{e_{s}},\mathbb{C}_{e_{t}})$ into different bins. The logarithmic numbers of the count values of each bin then constitute the interaction vector. This soft-interaction in a way is similar in the idea with the traditional link-based model Witten and Milne (2008), where the relatedness measure is based on the overlapping of in-coming links.
5.3 Attention-based CNN for temporal representation
For learning representation from entity temporal signals, the intuition is to model the low-level temporal correlation between two multivariate time series. Specifically, we learn to embed these time series of equal size $T$ into an Euclidean space, such that similar pairs are close to each other. Our embedding function takes the form of a convolutional neural network (CNN), shown in Figure 4. The architecture rests on four basic layers: a 1-D convolutional (that restricts the slide only along the time window dimension, following Zheng et al. (2014)), a batch-norm, an attention-based and a fully connected layer.
Convolution layer: A 1-D convolution operation involves applying a filter $\mathbf{w}_{f}\in\mathbf{R}^{1\times w\times D}$ (i.e., a matrix of weight parameters) to each subsequence $X^{i}_{e}$ of window size $m$ to produce a new abstraction.
$\mathit{q}_{i}=\mathbf{w}_{f}\mathsf{L^{i}}_{t:t+m-1,D}+b;\>\mathit{s}_{i}=BN(\mathit{q}_{i});\>\mathit{h}_{i}=ReLU(\mathit{s}_{i})$ (2)
where $\mathsf{L^{i}}_{t:t+w-1,D}$ denotes the concatenation of $w$ vectors in the lookup layer representing the subsequence $X^{i}_{e}$, $b$ is a bias term. The convolutional layer is followed by a batch normalization (BN) layer Ioffe and Szegedy (2015), to speed up the convergence and help improve generalization.
[…]
$\mathbb{F}^{h}_{i}[:,k]=A[k]\cdot\mathbb{F}^{c}_{i}[:,k],k=1\cdots T-w+1$ (3)
The time context vector $a$ is generated by a decay weight function, since each column $k$ in the vector is associated with a time $t_{k}$ which is $T-k+w$ time units away from studied time $t$.
Decay weight function: we leverage the Polynomial Curve for the function. $PD(t_{i},t)=\frac{1}{(t-t_{i})^{\alpha}+1}$, whereas $\alpha$ defines the decay rate. It is worth noting that when $\alpha$ is increased, the attention layer acts just like a pooling one  [^1] Note that, for clear visualization, we put flattening before attention layer in Figure 4. Stacking up multiple convolutional layers is possible, in this case $\lvert A\rvert$ is the size of the previous layer. The attention layer is only applied to the last convolution layer in our architecture. The output of the attention layer is then passed to a fully-connected layer with non-linear activation to obtain the temporal representation.
Figure 4: The attentional CNN for time series representation.
@@ 5.4 Learning and Optimization
Finally, we describe the optimization and training procedure of our network. We use a Logarithmic loss that can lead to better probability estimation at the cost of accuracy  [^2] Other ranking-based loss such as Hinge loss favours over sparsity and accuracy (in the sense of direct punishing misclassification via margins) at the cost of probability estimation. The logistic loss distinguishes better between examples whose supervision scores are close.. Our network minimizes the cross-entropy loss function as follows:
$L=-\frac{1}{N}\sum\limits_{i=1}^{N}[P_{\{e_{s},e_{1},e_{2}\}_{i}}\log\bar{y}_{i}\\ +(1-P_{\{e_{s},e_{1},e_{2}\}_{i}})\log(1-\bar{y}_{i})]+\lambda\lvert\theta\rvert_{2}^{2}$ (4)
where $N$ is the training size, $\bar{y}$ is the output of the sigmoid layer on the predicted label. $\theta$ contains all the parameters of the network and $\lambda\lvert\theta\rvert_{2}^{2}$ is the L2 regularization. $P_{\{e_{s},e_{(+)},e_{(-)}\}_{i}}$ is the probability that $e_{(+)}$ is ranked higher than $e_{(-)}$ derived from entity navigation, $P_{\{e_{s},e_{(+)},e_{(-)}\}_{i}}=y_{\{e_{s},e_{(+)}\}}^{t(i)}/(y_{\{e_{s},e_{(+)}\}}^{t(i)}+y_{\{e_{s},e_{(-)}\}}^{t(i)})$, where $t(i)$ is the observed time point of the training instance $i$. The network parameters are updated using Adam optimizer Kingma and Ba (2014).
@@ 6.3 Experimental Setup
Model Pearson ×$\times 100$ $\rho\times 100$ nDCG (proxy) nDCG (human)
@@
$Content_{Emb}$+$Graph_{Emb}$ 48.9 40.1 49.9 37.5 27.9 0.67 0.62 0.70 0.61 0.69 0.65
Base+$Content_{Emb}$ 67.1 54.2 53.4 43.7 26.5 0.67 0.69 0.71 0.61 0.72 0.74
Base+$Graph_{Emb}$ 55.2 50.2 41.3 31.5 35.5 0.71 0.75 0.78 ${}^{\mp}$ ${}^{\mp}$ ${}^{\mp}$
Trio 58.6 54.3 50.2 45.4 43.5 0.75 0.78 0.83 ${}^{\mp}$ ${}^{\mp}$ ${}^{\mp}$
@@ 6.7 Additional Analysis
Performances on different entity types. We demonstrate in Figures 4(b) and  4(c) the model performances on the person and event types. WLM performs poorer for the latter, that can be interpreted as link-based methods tend to slowly adapt for recent trending entities. The temporal models seem to capture these entites better.

### v5

@@ 1 Introduction
  • We formulate dynamic entity relatedness ranking and optimize directly for time-sensitive pairwise ordering rather than static similarity scores.
  • We introduce a temporal convolutional module with a monotonic time-decay weighting scheme, enabling the model to embed temporal signals and emphasize recency-dependent local patterns.
  • We propose a trio neural ranking framework that jointly models content-, graph-, and time-based views of entities, yielding jmutually informative representations tailored for ranking.
@@ 3.2 Problem Definition
We consider the task of quantifying and ranking the semantic relatedness between entities as it evolves over time. Rather than assuming a single static relatedness function, we allow the notion of relatedness to vary with time and model it as a family of functions $F=\{f_{t}\}_{t\in\mathcal{T}}$, where each $f_{t}$ reflects the relations observed at time $t$.
Dynamic Entity Relatedness.
For any time point $t$ and any pair of entities $(e_{s},e_{t})$ consisting of a source entity $e_{s}$ and a target entity $e_{t}$, the dynamic relatedness score is given by a function
$f_{t}:\mathcal{E}\times\mathcal{E}\rightarrow\mathbb{R}_{\geq 0}.$
We require each $f_{t}$ to satisfy the following properties:
[…]
We make no assumptions of symmetry, transitivity, or metric structure, allowing $f_{t}$ to capture complex and time-varying semantic interactions.
Dynamic Entity Relatedness Ranking.
Given a source entity $e_{s}$ and a time point $t$, the goal is to rank a set of candidate target entities $\{e_{1},\ldots,e_{n}\}$ according to their dynamic relatedness scores $f_{t}(e_{s},e_{1}),\ldots,f_{t}(e_{s},e_{n})$. The output is an ordering that reflects the relative strength of entity–entity associations at time $t$.
@@ 4.2 Neural Ranking Model Overview
Entity relatedness can in principle be modeled using a point-wise approach that directly predicts a scalar score for each entity pair. However, navigation data is highly skewed, and supervision from long-tail interactions is often noisy. Instead of learning a fully calibrated scoring function, we therefore adopt a pair-wise ranking strategy, which focuses on comparing candidate entities relative to one another. Pair-wise methods have the advantage of preserving partial orders of the underlying relatedness functions $f_{t}$ even when $f_{t}$ is not globally transitive Cheng et al. (2012), making them well-suited for dynamic entity ranking.
Our architecture follows an interaction-based design in which the model directly processes the triplet $(e_{s},e_{(+)},e_{(-)})$ to learn their relative compatibility. In contrast to Siamese or representation-based models Chopra et al. (2005), our networks do not share parameters across branches. Parameter sharing would implicitly enforce symmetry $f_{t}(e_{s},e)=f_{t}(e,e_{s})$, contradicting the inherently asymmetric nature of the relatedness function (Section 3.2).
Each branch of the model is a feed-forward network $\psi$ with input $z_{0}$, hidden layers $z_{1},\ldots,z_{n-1}$, and output $z_{n}$, where
$z_{i}=\sigma(\mathbf{W}_{i}z_{i-1}+\mathbf{b}_{i}),$
and $\sigma$ is a nonlinear activation such as ReLU. Under the trio setup, the overall pair-wise score is the sum of the outputs from the temporal, graph, and content networks:
$\phi(e_{s},e_{(+)},e_{(-)})=\phi_{\mathrm{time}}+\phi_{\mathrm{graph}}+\phi_{\mathrm{content}}.$
The next section details the input representations $z_{0}$ used by each network.
@@ 5.1 Content-based representation learning
To obtain content representations, we leverage both the textual document (word-based) and the link profile (entity-based) of each entity, as introduced in Section 3.1. Given the large word and entity vocabularies, we employ the word hashing technique of Huang et al. (2013), which maps each token to a set of character trigrams. Let $\mathcal{V}$ denote the trigram vocabulary and $\mathsf{E}:\mathcal{V}\!\rightarrow\!\mathbb{R}^{m}$ the embedding function. We also learn a global importance weight $\mathsf{w}:\mathcal{V}\!\rightarrow\!\mathbb{R}_{\geq 0}$.
For an entity $e$, let $e_{w}=(v_{1},\ldots,v_{n_{e}})$ denote the sequence of hashed trigrams extracted from its document. We represent $e$ by a weighted compositional embedding:
$\phi_{\mathrm{word}}(e)=\sum_{j=1}^{n_{e}}\mathsf{w}(v_{j})\,\mathsf{E}(v_{j})\;\in\mathbb{R}^{m}.$
This formulation corresponds to a linear bag-of-subword encoder and is equivalent to the expected embedding of a token drawn proportionally to its trigram weights.
For the entity-based representation, we analogously define $e_{\mathrm{ent}}=(u_{1},\ldots,u_{k_{e}})$ as the hashed tokens obtained from the surface forms of entities linked from $e$, and compute:
$\phi_{\mathrm{ent}}(e)=\sum_{j=1}^{k_{e}}\mathsf{w}(u_{j})\,\mathsf{E}(u_{j}).$
The final content representation is their concatenation:
$\phi_{\mathrm{content}}(e)=[\phi_{\mathrm{word}}(e);\phi_{\mathrm{ent}}(e)]\in\mathbb{R}^{2m}.$
For a triplet $(e_{s},e_{(+)},e_{(-)})$, the concatenated content representations serve as the input to the content sub-network.
@@ 5.2 Graph-based representation
For graph-based representations, we follow the DeepWalk framework Perozzi et al. (2014). Let $G=(V,E)$ be the entity graph, where $V$ is the set of entities and edges correspond to hyperlinks. DeepWalk generates random walks $\mathbb{S}_{e}=(v_{1},\ldots,v_{L})$ from each entity $e$ and optimizes a Skip-gram objective:
$\max_{\mathsf{G}}\sum_{e\in V}\sum_{v_{j}\in\mathbb{S}_{e}}\sum_{v_{k}\in\mathcal{N}(v_{j})}\log\mathbb{P}\big(v_{k}\mid\mathsf{G}(v_{j})\big),$
where $\mathsf{G}:V\!\rightarrow\!\mathbb{R}^{d}$ is the learned graph embedding and $\mathcal{N}(v_{j})$ is the context window. This yields graph-aware embeddings capturing co-occurrence structure under random walks.
Given two entities $e_{s}$ and $e_{t}$, let
$\mathbb{C}_{e_{s}}=\{\mathsf{G}(v):v\in\mathbb{S}_{e_{s}}\},\qquad\mathbb{C}_{e_{t}}=\{\mathsf{G}(u):u\in\mathbb{S}_{e_{t}}\}$
denote their graph-context embeddings. For every pair $(x,y)\in(\mathbb{C}_{e_{s}},\mathbb{C}_{e_{t}})$, we compute the cosine similarity:
$\mathrm{sim}(x,y)=\frac{\langle x,y\rangle}{\lVert x\rVert\,\lVert y\rVert}.$
Following Guo et al. (2016), we discretize these similarities into $B$ fixed histogram bins. Let $H_{b}$ be the count of similarities falling into bin $b$. The interaction vector is:
$\phi_{\mathrm{graph}}(e_{s},e_{t})=\big[\log(1+H_{1}),\;\ldots,\;\log(1+H_{B})\big]\in\mathbb{R}^{B}.$
This histogram encodes soft matching of graph neighborhoods, analogous to classical link-based relatedness measures Witten and Milne (2008), but operating at the embedding level.
5.3 Temporal Representation via Time-Weighted Convolution
To learn representations from temporal signals, we treat each entity’s multivariate time series as a sequence of $T$ aligned observations. Our goal is to embed such sequences into a Euclidean space in which temporally similar entities lie close together. We employ a 1-D convolutional architecture that extracts local temporal patterns and then reweights these patterns according to their temporal proximity to the prediction time. The architecture consists of: (i) a temporal convolution layer, (ii) batch normalization, (iii) a time-decay weighting mechanism, and (iv) a fully connected projection layer.
Temporal convolution.
Let $X=(x_{1},\ldots,x_{T})$ denote an entity’s time series, where $x_{t}\in\mathbb{R}^{D}$ is a $D$-dimensional feature vector at time $t$ (e.g., a scalar popularity signal or a small feature bundle). A 1-D convolution with kernel width $w$ applies a filter $\mathbf{W}\in\mathbb{R}^{w\times D}$ and bias $b\in\mathbb{R}$ to each contiguous window $(x_{t-w+1},\ldots,x_{t})$:
$q_{t}=\langle\mathbf{W},X_{t-w+1:t}\rangle+b,\qquad t=w,\ldots,T.$
Batch normalization and a ReLU activation yield
$h_{t}=\mathrm{ReLU}\!\left(\mathrm{BN}(q_{t})\right),\qquad t=w,\ldots,T.$
The resulting sequence $\mathbf{h}=(h_{w},h_{w+1},\ldots,h_{T})$ captures local temporal patterns (bursts, surges, and short-term trends).
Time-decay weighting.
Not all temporal patterns should contribute equally: patterns closer to the prediction time are intuitively more relevant. We therefore introduce a deterministic, strictly positive decay function $A_{t}$ that assigns higher weight to features extracted from recent time steps.
For the convolution output index $t$, denote its temporal distance from the prediction point by $\Delta_{t}=T-t$. We define:
$A_{t}=\frac{1}{(1+\Delta_{t})^{\alpha}},\qquad\alpha>0,\quad t=w,\ldots,T.$
The weighted convolutional representation is then:
$\tilde{h}_{t}=A_{t}\cdot h_{t},\qquad t=w,\ldots,T.$
This mechanism is a deterministic “soft focusing” operation: unlike learned attention, the weights are fixed by temporal proximity, providing an interpretable, monotonic recency bias.
Projection.
Finally, the weighted sequence is flattened and passed through a nonlinear fully connected layer:
$z=\sigma\!\left(\mathbf{W}_{\mathrm{fc}}\,[\tilde{h}_{w};\tilde{h}_{w+1};\ldots;\tilde{h}_{T}]+\mathbf{b}_{\mathrm{fc}}\right),$
which yields the final temporal embedding $z\in\mathbb{R}^{d}$. Only the last convolutional layer is time-weighted, though deeper stacks of convolutional layers can be used.
This design cleanly separates (i) local temporal pattern extraction via convolution, and (ii) global temporal relevance modulation via a principled, monotonic decay, yielding an interpretable and robust temporal representation.
Figure 4: Architecture of the temporal convolution module. Local temporal patterns are extracted by a 1-D convolution and then modulated by a monotonic time-decay weighting before projection into the final temporal embedding space.
@@ 5.4 Learning and Optimization
Our model learns a pairwise preference function over entity candidates. For each training instance $i$ consisting of a source entity $e_{s}$ and a positive–negative pair $(e_{(+)},e_{(-)})$, the model produces a score $f_{t(i)}(e_{s},e)$ for each candidate. The probability that $e_{(+)}$ should be ranked above $e_{(-)}$ is defined through a logistic link applied to the score difference:
$\bar{y}_{i}=\sigma\!\big(f_{t(i)}(e_{s},e_{(+)})-f_{t(i)}(e_{s},e_{(-)})\big).$
To obtain supervision, we construct soft preference targets from navigation statistics:
$\tilde{P}_{i}=\frac{y^{t(i)}_{\{e_{s},e_{(+)}\}}}{y^{t(i)}_{\{e_{s},e_{(+)}\}}+y^{t(i)}_{\{e_{s},e_{(-)}\}}},$
which provide a smoothed estimate of how often users prefer $e_{(+)}$ over $e_{(-)}$ when navigating from $e_{s}$ at time $t(i)$.
We train the network using the regularized cross-entropy objective:
$L=-\frac{1}{N}\sum_{i=1}^{N}\Big(\tilde{P}_{i}\log\bar{y}_{i}+(1-\tilde{P}_{i})\log(1-\bar{y}_{i})\Big)+\lambda\lVert\theta\rVert_{2}^{2},$
where $\theta$ denotes all model parameters. This loss is a smooth pairwise surrogate for ranking and provides calibrated probability estimates, enabling the model to distinguish fine differences in temporal relatedness. Parameters are optimized using Adam Kingma and Ba (2014).
@@ 6.3 Experimental Setup
Model Pearson $\times 100$ $\rho\times 100$ nDCG (proxy) nDCG (human)
@@
$Content_{Emb}$$Graph_{Emb}$ 48.9 40.1 49.9 37.5 27.9 0.67 0.62 0.70 0.61 0.69 0.65
[…]
Trio 58.6 54.3 50.2 45.4 43.5 0.75 0.78 0.83 0.74 0.82 0.85
@@ 6.7 Additional Analysis
Performances on different entity types. We demonstrate in Figures 5(b) and  5(c) the model performances on the person and event types. WLM performs poorer for the latter, that can be interpreted as link-based methods tend to slowly adapt for recent trending entities. The temporal models seem to capture these entites better.